# ViewFormer: Exploring Spatiotemporal Modeling for Multi-View 3D Occupancy Perception via View-Guided Transformers

### v1

@@ Abstract
3D occupancy, an advanced perception technology for driving scenarios, represents the entire scene without distinguishing between foreground and background by quantifying the physical space into a grid map. The widely adopted projection-first deformable attention, efficient in transforming image features into 3D representations, encounters challenges in aggregating multi-view features due to sensor deployment constraints. To address this issue, we propose our learning-first view attention mechanism for effective multi-view feature aggregation. Moreover, we showcase the scalability of our view attention across diverse multi-view 3D tasks, such as map construction and 3D object detection. Leveraging the proposed view attention as well as an additional multi-frame streaming temporal attention, we introduce ViewFormer, a vision-centric transformer-based framework for spatiotemporal feature aggregation. To further explore occupancy-level flow representation, we present FlowOcc3D, a benchmark built on top of existing high-quality datasets. Qualitative and quantitative analyses on this benchmark reveal the potential to represent fine-grained dynamic scenes. Extensive experiments show that our approach significantly outperforms prior state-of-the-art methods. The codes and benchmark will be released soon.
@@
3D Occupancy Occupancy flow Multi-view Interaction Temporal modeling Spatiotemporal Modeling Streaming video pipeline
@@ 1 Introduction
The vision-centric autonomous driving (AD) systems are attracting extensive attention in recent years with remarkable cost-effectiveness, promoting the research domain focused on perceiving the real 3D world from 2D images. 3D object detection is a traditional task in an AD system, representing foreground objects with limited categories. As illustrated in Fig. 1(a), the pedestrian is detected as a 3D bounding box, while the suitcase, which is uncommon in driving scenarios, is hardly defined. Once such a suitcase appears on the road, an AD system relying only on object detection tasks is unable to secure the driving safety. In contrast, the 3D occupancy representation, unifying the concept of foreground and background, and quantifying the entire 3D space into voxel-wise
@@
cells with semantic labels, shows superior performance, where objects like suitcases can be effectively defined as category-agnostic occupancies as shown in Fig. 1(b). Beyond static occupancy, occupancy flow is crucial for representing dynamic scenes as shown in Fig. 1(d). In this paper, we introduce ViewFormer, a transformer-based framework designed to predict 3D occupancy and occupancy flow with multi-camera images as input.
@@
To gain a better understanding, let’s review the projection-first method illustrated in Fig. 2(a), which projects the 3D reference point of a voxel query onto images first and then performs deformable attention [42]. However, a critical limitation arises when a 3D reference point, fixed during training, cannot be projected onto a specific camera due to the limited image size. As a consequence, features from this camera are masked out, leading to an inability to collect features from this camera throughout the entire dataset. Notably, this issue is widespread, as seen in scenarios like nuScenes [5, 9], where numerous 3D reference points can only be projected onto a single camera due to sensor deployment. Additionally, as the deformable attention utilized in the method learns sample points on the image plane, the 2D sample area corresponding to a 3D object undergoes rapid variations with changes in depth due to perspective transformation, such scale inconsistency poses a convergence challenge to 3D perception tasks.
@@ 2 Related Work
3D Scene Reconstruction. The occupancy network, introduced by Tesla [1], brings the concept of the occupancy grid map, which has long been utilized in the domains of robotic mapping and planning [30, 8], into the AD field. Semantic scene completion [32] is similar to the goals of occupancy tasks discussed in this paper. MonoScene [6] leverages a U-Net architecture to infer dense 3D occupancy from a single image. VoxFormer [16] introduces depth estimation to guide voxel queries, and OccDepth [25] adopts stereo depth information to improve occupancy prediction. TPVFormer [13] proposes a tri-perspective view representation to aggregate image features. FB-OCC [19] constructs 3D features via forward-backward transformations. OccNet [31] introduces a cascade voxel decoder to achieve 3D occupancy prediction. In addition, [31, 38] contribute high-quality occupancy benchmarks to facilitate the research community. Although [31] assigns coarse object-level velocity directly to object occupancies, fine-grained occupancy-level flow representation still remain unexplored, by which we are motivated to present our occupancy-level flow dataset FlowOcc3D, as well as a feasible framework based on our view attention and streaming temporal attention for motion 3D occupancy prediction.
@@ 5.1 Datasets
Table 3: 3D Occupancy and Occupancy Flow Prediction on FlowOcc3D benchmark. Our ViewFormer achieves the best result among previous SOTAs. *: we add a flow head for flow prediction to BEVFormer [18] and FB-OCC [19] and retrain them on FlowOcc3D.
@@ 5.2 Main Results
3D Occupancy on Occ3D. We compare our ViewFormer with previous state-of-the-art methods on the 3D occupancy task in Tab. 1, where the baselines BEVFormer [18], TPVFormer [13], CTF-Occ [40] and FB-OCC [19] adopt the projection-first spatial interaction method as analyzed in Sec. 1. In terms of training setup, we follow FB-OCC [19], the winner of 3D occupancy challenge at CVPR 2023 [34, 31], which includes 90-epoch training, the same pretrained weights, and depth supervision for image backbone. As shown in Tab. 1, in contrast to the strong baseline FB-OCC [19], which utilizes longer video sequences, our model achieves 2.74 mIoU improvement, proving the superiority of our ViewFormer.
3D Occupancy on OpenOcc. We conduct comparison experiments on the OpenOcc benchmark in Tab. 2, where we adopt the same configuration as OccNet [31], with an image size of 450 $\times$ 800. Our approach demonstrates significant performance improvements over the baselines. Especially, we surpass the state-of-the-art OccNet [31] by 4.36 on mIoU and 2.71 on $\text{IoU}_{geo}$ when utilizing Res50 as the backbone.
@@ 5.3 Ablations and Analysis
5.3.1 View Attention.
To validate the effectiveness of our view attention, we conduct experiments as presented in Tab. 4. Since this module serves as a spatial interaction and is not involved in temporal modeling, we temporarily disable the flow head. The findings are as below. 1) We replace our view attention module with the projection-first deformable attention as BEVformer [18], the results show that our learning-first view attention leads to an improvement of 1.22 mIoU and 0.43 IoUgeo under approximate computational complexity, revealing the limitation of the commonly used projection-first method for multi-view feature aggregation. 2) Learning sample points directly in the ego-centric perception coordinate system instead of the query-specific view coordinate (VC) system results in performance degradation of -0.75 mIoU, which validates our motivation mentioned in Sec. 1 that the rotation invariance introduced by our view attention effectively reduces learning complexity, accelerates convergence, and enhances accuracy. Detailed convergence experiments can be found in supplementary materials.
@@
5.3.2 Visualization Analysis.
We visualize our view attention in Fig. 5. The green cross mark denotes a query’s reference point, which can only be projected onto the left image. The learned 3D points are projected onto multi-view images represented by filled circles, with colors to indicate attention weights. In such scenarios, the projection-first method, like BEVFormer [18] that only gathers features from the image where the reference point can be projected, fails to collect features from the right image, as discussed in Sec. 1. In contrast, our method is competent to gather features from both images, allowing for constructing more semantically informative and robust features for this query.
@@
5.3.3 Application of View Attention
We also apply our view attention to other tasks to assess its scalability, including MapTR [20] for HD map construction and DETR3D [39] for 3D object detection, both of them collect multi-view image features in a similar way to the projection-first method analyzed in Sec. 1. We replace their corresponding feature collection module with our view attention. The evaluation metrics are available in their respective literatures. We train MapTR for 24 epochs with an image size of $324\times 576$, and Res50 as the backbone. Tab. 5 indicates our ViewAttn-MapTR achieves a substantial 5.82% improvement on mAP, demonstrating a remarkable increase in convergence speed. For training DETR3D, we use Res101-DCN as the backbone without CBGS [41]. Tab. 6 shows our ViewAttn-DETR3D also results in 4.1% and 1.9% improvements on mAP and NDS respectively. Unlike a considerable amount of recent researches mainly focus on temporal modeling, lacking the study of the fundamental spatial interaction for multi-camera 3D perception, our work reveals the limitations of the widely used projection-first method and highlights substantial research opportunities that remain unexplored.
@@ 5.4 Qualitative Results
Figure 6: Qualitative results of 3D occupancy and occupancy flow prediction. 3D occupancies are color-coded according to semantic categories. For flow visualization, the color and brightness represent the motion direction and magnitude respectively.
We present qualitative results of our ViewFormer in Fig. 6. For the occupancy task, 3D occupancies are color-coded according to semantic categories, and to visualize the flow, we utilize color and brightness to represent the flow direction
Figure 7: Visualization on flow predictions. Compared with the flow predictions (a) supervised by object-level flow, the flow predictions (b) supervised by our occupancy-level FlowOcc3D show more reasonable results for a turning car.
and magnitude respectively, following the convention of the optical flow fields [33, 2]. An additional animated video is also provided in the supplementary material. Moreover, we demonstrate a qualitative comparison for flow predictions supervised by object-level flow and occupancy-level flow in Fig. 7, the model trained with our occupancy-level FlowOcc3D dataset achieves more reasonable results for a turning car, providing fine-grained motion information for AD systems.
@@ 6 Conclusion
Limitations and Future Work. We generate occupancy flow only based on foreground object annotations, background motion information is lacking. Incorporating data with complete motion information from real dynamic scenes would be a valuable avenue for future developments in AD systems.

### v2

@@ Abstract
3D occupancy, an advanced perception technology for driving scenarios, represents the entire scene without distinguishing between foreground and background by quantifying the physical space into a grid map. The widely adopted projection-first deformable attention, efficient in transforming image features into 3D representations, encounters challenges in aggregating multi-view features due to sensor deployment constraints. To address this issue, we propose our learning-first view attention mechanism for effective multi-view feature aggregation. Moreover, we showcase the scalability of our view attention across diverse multi-view 3D tasks, including map construction and 3D object detection. Leveraging the proposed view attention as well as an additional multi-frame streaming temporal attention, we introduce ViewFormer, a vision-centric transformer-based framework for spatiotemporal feature aggregation. To further explore occupancy-level flow representation, we present FlowOcc3D, a benchmark built on top of existing high-quality datasets. Qualitative and quantitative analyses on this benchmark reveal the potential to represent fine-grained dynamic scenes. Extensive experiments show that our approach significantly outperforms prior state-of-the-art methods. The codes are available at https://github.com/ViewFormerOcc/ViewFormer-Occ.
@@
3D Occupancy Occupancy flow Multi-view Interaction Spatiotemporal Modeling Streaming video pipeline
@@ 1 Introduction
The vision-centric autonomous driving (AD) systems are attracting extensive attention in recent years, promoting the research domain perceiving the real 3D world from 2D images. 3D object detection is a traditional task, representing foreground objects with limited categories. As illustrated in Fig. 1(a), the pedestrian is detected as a 3D bounding box, while the uncommon suitcase is hardly defined in driving scenarios. Once such a suitcase appears on the road, an AD system relying only on object detection is unable to secure the driving safety. In contrast, the 3D occupancy representation, unifying the concept of foreground and background, and quantifying the entire 3D space into voxel-wise cells with
@@
semantic labels, shows superior performance, where objects like suitcases can be effectively defined as category-agnostic occupancies as shown in Fig. 1(b). Beyond static occupancy, occupancy flow is crucial for representing dynamic scenes as shown in Fig. 1(d). In this paper, we introduce ViewFormer, a transformer-based framework designed to predict 3D occupancy and occupancy flow with multi-camera images as input.
@@
To gain a better understanding, let’s review the projection-first method illustrated in Fig. 2(a), which projects the 3D reference point of a voxel query onto images first and then performs deformable attention [42]. However, a critical limitation arises when a 3D reference point, fixed during training, is projected outside the image size for a specific camera, the projection-first method no longer applies deformable attention to extracting features for this reference point. As a consequence, features from this camera are masked out throughout the entire dataset. Notably, this issue is widespread, as seen in scenarios like nuScenes [5, 9], where numerous 3D reference points can only be projected onto a single camera due to sensor deployment. Additionally, as the deformable attention utilized in the method learns sample points on the image plane, the 2D sample area corresponding to a 3D object undergoes rapid variations with changes in depth due to perspective transformation, such scale inconsistency poses a convergence challenge to 3D perception tasks.
@@ 2 Related Work
3D Scene Reconstruction. The occupancy network, introduced by Tesla [1], brings the concept of the occupancy grid map, which has long been utilized in the domains of robotic mapping and planning [30, 8], into the AD field. Semantic scene completion [32] is similar to the goals of occupancy tasks discussed in this paper. MonoScene [6] leverages a U-Net architecture to infer dense 3D occupancy from a single image. VoxFormer [16] introduces depth estimation to guide voxel queries, and OccDepth [25] adopts stereo depth information to improve occupancy prediction. TPVFormer [13] proposes a tri-perspective view representation to aggregate image features. FB-OCC [19] constructs 3D features via forward-backward transformations. In addition, [31, 38] contribute high-quality occupancy benchmarks to facilitate the research community. Although [31] assigns coarse object-level velocity directly to object occupancies, fine-grained occupancy-level flow representation still remain unexplored, by which we are motivated to present our occupancy-level flow dataset FlowOcc3D, as well as a feasible framework based on our view attention and streaming temporal attention for motion 3D occupancy prediction.
@@ 5.1 Datasets
Table 3: 3D Occupancy and Occupancy Flow Prediction on FlowOcc3D. Our ViewFormer achieves the best result among previous SOTAs. *: we add a flow head for flow prediction to BEVFormer [18] and FB-OCC [19] and retrain them on FlowOcc3D.
@@ 5.2 Main Results
3D Occupancy on Occ3D. We compare our ViewFormer with previous state-of-the-art methods on the 3D occupancy task in Tab. 1, where the baselines BEVFormer [18], TPVFormer [13], CTF-Occ [40] and FB-OCC [19] adopt the projection-first spatial interaction method as analyzed in Sec. 1. In terms of training setup, we follow FB-OCC [19], the winner of 3D occupancy challenge of CVPR 2023 [34, 31], which includes 90-epoch training, the same pretrained weights, and depth supervision for image backbone. As shown in Tab. 1, in contrast to the strong baseline FB-OCC [19], which utilizes longer video sequences, our model achieves 2.74 mIoU improvement, proving its superiority.
3D Occupancy on OpenOcc. We conduct experiments on OpenOcc in Tab. 2, where we adopt the same configuration as OccNet [31], with an image size of 450 $\times$ 800. Our approach demonstrates significant performance improvements over the baselines. Especially, we surpass the state-of-the-art OccNet [31] by 4.36 on mIoU and 2.71 on $\text{IoU}_{geo}$ when utilizing Res50 as the backbone.
@@ 5.3 Ablations and Analysis
View Attention. To validate the effectiveness of our view attention, we conduct experiments as presented in Tab. 4. Since this module serves as a spatial interaction and is not involved in temporal modeling, we temporarily disable the flow head. The findings are as below. 1) We replace our view attention module with the projection-first deformable attention as BEVformer [18], the results show that our learning-first view attention leads to an improvement of 1.22 mIoU and 0.43 IoUgeo under approximate computational complexity, revealing the limitation of the commonly used projection-first method for multi-view feature aggregation. 2) Learning sample points directly in the ego-centric perception coordinate system instead of the query-specific view coordinate (VC) system results in performance degradation of -0.75 mIoU, which validates our motivation mentioned in Sec. 1 that the rotation invariance introduced by our view attention effectively reduces learning complexity, accelerates convergence, and enhances accuracy. Detailed convergence experiments can be found in supplementary materials.
@@
Visualization Analysis. We visualize our view attention in Fig. 5. The green cross mark denotes a query’s reference point, which can only be projected onto the left image. The learned 3D points are projected onto multi-view images represented by filled circles, with colors to indicate attention weights. In such scenarios, the projection-first method, like BEVFormer [18] that only gathers features from the image where the reference point can be projected, fails to collect features from the right image, as discussed in Sec. 1. In contrast, our method is competent to gather features from both images, allowing for constructing more semantically informative and robust features for this query.
@@
Application of View Attention. We also apply our view attention to other tasks to assess its scalability, including MapTR [20] for HD map construction and DETR3D [39] for 3D object detection, both of them collect multi-view image features in a similar way to the projection-first method analyzed in Sec. 1. We replace their corresponding feature collection module with our view attention. The evaluation metrics are available in their respective literatures. We train MapTR for 24 epochs with an image size of $324\times 576$, and Res50 as the backbone. Tab. 5 indicates our ViewAttn-MapTR achieves a substantial 5.82% improvement on mAP, demonstrating a remarkable increase in convergence speed. For training DETR3D, we use Res101-DCN as the backbone without CBGS [41]. Tab. 6 shows our ViewAttn-DETR3D also results in 4.1% and 1.9% improvements on mAP and NDS respectively. Unlike a considerable amount of recent researches mainly focusing on temporal modeling, and lacking the study of the fundamental spatial interaction for multi-camera 3D perception, our work reveals the limitations of the widely used projection-first method and highlights substantial research opportunities that remain unexplored.
@@ 5.4 Qualitative Results
Figure 6: Qualitative results of 3D occupancy and occupancy flow prediction. 3D occupancies are color-coded according to semantic categories. For flow visualization, we utilize color and brightness to represent the flow direction and magnitude respectively, following the convention of the optical flow fields [33, 2].
Figure 7: Visualization on flow predictions. (a) Flow predictions supervised by object-level flow. (b) Flow predictions supervised by our FlowOcc3D.
We present qualitative results of our ViewFormer in Fig. 6. An additional animated video is also provided in the supplementary material. Moreover, we demonstrate a qualitative comparison for flow predictions supervised by object-level flow (a) and occupancy-level flow (b) in Fig. 7, the model trained with our occupancy-level FlowOcc3D dataset achieves more reasonable results for a turning car, providing fine-grained motion information for AD systems.